# Sparse Subnetwork Enhancement for Underrepresented Languages in Large Language Models

### v1

@@ Acknowledgments
Appendix A Language Neuron Overlap and Individual Distributions
Figure 8: Language neuron overlap for all 12 languages for Llama-3.1-8B.
Figure 9: Language neuron overlap for all 12 languages for Mistral-Nemo-12B.
Figure 10: Neuron distributions over the layers of Mistral-Nemo-12B for all 12 languages.
Lang. # Neurons # Trainable Param-s %
@@ Appendix A Language Neuron Overlap and Individual Distributions
Figure 12: Language neuron distributions for all 12 languages for Llama-3.1-8B-12B.
Figure 13: Language neuron distributions for all 12 languages for Mistral-Nemo-12B.
Appendix B Validation Losses
(a) Afrikaans validation loss curves for Llama-3.1-8B.
(b) Welsh validation loss curves for Llama-3.1-8B.
(c) Estonian validation loss curves for Llama-3.1-8B.
(d) Icelandic validation loss curves for Llama-3.1-8B.
(e) Georgian validation loss curves for Llama-3.1-8B.
(f) Lithuanian validation loss curves for Llama-3.1-8B.
(a) Latvian validation loss curves for Llama-3.1-8B.
(b) Macedonian validation loss curves for Llama-3.1-8B.
(c) Maltese validation loss curves for Llama-3.1-8B.
(d) Nepali validation loss curves for Llama-3.1-8B.
(e) Slovak validation loss curves for Llama-3.1-8B.
(f) Slovenian validation loss curves for Llama-3.1-8B.
Appendix C Downstream Results
Metric Base Full MLP LoRA Random Target
@@ Appendix C Downstream Results
arc challenge 70.90 25.75 35.45 64.55 62.21 71.24
arc easy 85.79 26.14 53.33 79.12 81.75 85.26
hellaswag 40.38 14.68 26.13 35.22 36.04 38.72
piqa 71.87 39.17 50.00 73.07 72.03 71.71
winogrande 51.38 49.17 50.43 50.91 47.04 51.46
mmlu 58.07 27.63 37.30 51.60 54.74 56.43
Avg. 63.07 30.42 42.11 59.08 58.97 62.47
Table 3: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 16.05 22.74 56.52 54.52 68.23
arc easy 85.79 16.84 30.70 74.56 76.14 84.04
hellaswag 40.38 12.31 25.27 29.93 30.88 40.79
piqa 71.87 3.97 49.46 58.98 66.97 68.39
winogrande 51.62 20.28 49.57 49.17 51.46 50.51
mmlu 58.07 22.27 25.67 45.79 48.14 56.69
Avg. 63.10 15.29 33.90 52.49 54.68 61.44
Table 4: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 22.07 36.45 58.86 63.55 72.58
arc easy 85.79 21.40 55.09 77.89 81.58 84.74
hellaswag 40.38 13.92 22.14 36.11 37.06 41.82
piqa 71.87 40.10 54.35 59.63 72.96 75.24
winogrande 51.38 41.83 50.20 51.14 52.01 53.83
mmlu 58.07 20.64 33.90 48.66 52.84 57.35
Avg. 63.07 26.66 42.02 55.38 60.00 64.26
Table 5: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 21.74 37.79 56.19 59.20 74.25
arc easy 85.79 23.33 55.96 74.39 76.84 85.61
hellaswag 40.38 8.29 25.24 32.65 27.75 44.67
piqa 71.87 26.55 54.79 69.48 61.48 77.31
winogrande 51.38 40.02 50.28 52.01 52.17 54.54
mmlu 58.07 17.18 35.86 46.77 45.40 58.92
Avg. 63.07 22.85 43.32 55.25 53.81 65.88
Table 6: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 9.03 10.70 71.57 63.55 69.90
arc easy 85.79 10.18 15.44 85.26 82.98 84.56
hellaswag 40.38 1.24 6.33 42.39 33.13 41.56
piqa 71.87 12.13 8.81 73.67 64.53 71.16
winogrande 51.38 3.95 37.57 53.12 50.91 51.38
mmlu 58.07 4.51 18.42 58.00 55.26 56.69
Avg. 63.07 6.84 16.21 64.00 58.39 62.54
Table 7: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 20.74 37.79 54.52 67.89 70.23
arc easy 85.79 27.89 59.12 74.04 80.53 85.26
hellaswag 40.38 20.49 29.02 31.45 38.14 41.37
piqa 71.87 39.28 51.63 57.56 69.15 74.21
winogrande 51.38 51.07 49.96 50.43 50.36 53.35
mmlu 58.07 27.24 36.64 45.85 52.19 57.02
Avg. 63.07 31.12 44.03 52.31 59.71 63.58
Table 8: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 20.07 25.42 53.51 68.23 69.23
arc easy 85.79 27.54 36.67 70.70 81.58 84.56
hellaswag 40.38 14.48 17.34 31.23 35.03 41.73
piqa 71.87 39.50 43.69 55.93 71.82 74.59
winogrande 51.62 47.43 46.80 50.04 50.43 55.17
mmlu 58.07 24.89 30.70 45.59 53.43 59.63
Avg. 63.10 28.98 33.44 51.17 60.09 64.15
Table 9: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 31.10 43.48 69.90 66.22 69.23
arc easy 85.79 39.82 64.39 84.39 82.11 85.61
hellaswag 40.38 26.41 25.06 37.91 32.57 36.43
piqa 71.87 49.67 53.26 73.72 68.44 67.36
winogrande 51.62 49.41 49.25 52.96 51.78 52.09
mmlu 58.07 30.96 39.65 56.69 54.67 57.48
Avg. 63.10 37.90 45.85 62.60 59.30 61.37
Table 10: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 22.07 38.13 49.16 62.21 71.57
arc easy 85.79 21.23 56.84 72.46 82.11 83.86
hellaswag 40.38 9.70 26.72 31.45 35.07 39.00
piqa 71.87 42.11 50.87 61.97 72.85 73.34
winogrande 51.54 51.22 49.17 49.96 53.67 50.91
mmlu 58.07 26.26 32.66 45.00 52.58 57.87
Avg. 63.09 28.77 42.40 51.67 59.75 62.76
Table 11: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 12.71 35.45 65.89 67.56 71.57
arc easy 85.79 19.47 49.12 81.75 82.98 84.39
hellaswag 40.38 16.38 14.32 37.46 36.68 36.31
piqa 71.87 36.83 51.25 62.35 73.83 71.11
winogrande 51.38 13.58 47.04 50.12 54.06 50.99
mmlu 58.07 17.50 33.18 54.08 54.08 57.74
Avg. 63.07 19.41 38.39 58.61 61.53 62.02
Table 12: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 21.74 36.45 61.20 62.21 71.57
arc easy 85.79 23.86 51.05 78.25 78.60 85.44
hellaswag 40.38 15.04 27.87 40.04 36.74 43.68
piqa 71.87 43.25 54.30 59.25 70.29 71.76
winogrande 51.54 38.91 50.43 51.07 50.91 51.70
mmlu 58.07 21.36 35.99 52.25 51.53 57.67
Avg. 63.09 27.36 42.68 57.01 58.38 63.64
Table 13: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
@@
arc challenge 70.90 21.74 34.78 53.51 67.56 67.89
arc easy 85.79 28.60 36.49 71.23 82.46 83.51
hellaswag 40.38 8.79 24.92 34.08 39.72 36.48
piqa 71.87 45.65 56.15 70.13 72.03 75.08
winogrande 51.46 49.88 50.67 49.64 50.43 52.25
mmlu 58.07 27.43 29.13 43.31 55.06 56.69
Avg. 63.08 30.35 38.69 53.65 61.21 61.98
Table 14: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
C.1 Mistral-Nemo-12B
@@
Table 15: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 16: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 17: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 18: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 19: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 20: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 21: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 22: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 23: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 24: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 25: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak. Best performing configurations are highlighted in dark green, second-best in light green.
@@
Table 26: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian. Best performing configurations are highlighted in dark green, second-best in light green.
C.2 Generation Prompts
Task Prompts

### v2

@@ Acknowledgments
Appendix A Limitations
While our results show that sparse, language-specific subnetwork fine-tuning is effective across multiple models and languages, several limitations remain.
First, the effectiveness of targeted subnetwork adaptation depends on accurately identifying language-specific neurons. This process can be sensitive to the quality, coverage, and representativeness of the monolingual data used for neuron identification, particularly for languages with limited or noisy corpora.
Second, our experiments primarily focus on mid- and low-resource languages with sufficient post-training data. How well the method scales to extremely low-resource settings or to languages that are highly distant from those seen during pre-training remains an open question.
Third, we adopt a fixed selection threshold and do not systematically vary the capacity of language-specific subnetworks (i.e., the parameter $\rho$). Our results suggest that different languages may require different amounts of capacity to effectively capture language-specific knowledge, and that the interaction between subnetwork size and available training data warrants further study, potentially at the per-language or per-language-family level.
Finally, due to computational constraints, we did not run Full fine-tuning, FFN-only, LoRA, or IA3 baselines for Mistral-Nemo-12B and Aya-Expanse-8B. Based on our observations with Llama-3.1-8B, we expect these dense and PEFT methods to exhibit similar behavior on these models, but formal verification remains future work.
Appendix B Language Neuron Details
B.1 Language Subnetworks
Figures B.1 and 10 illustrate how language-specific neurons are distributed across layers for each model. While the absolute number of selected neurons varies substantially across languages and architectures, consistent structural patterns emerge. In particular, subnetworks tend to be sparse in early layers and increasingly concentrated in later layers, suggesting that language-specific specialization primarily occurs at higher levels of abstraction. Differences in subnetwork size across languages likely reflect a combination of pre-training exposure, typological similarity to high-resource languages, and model-specific capacity allocation.
Figure 7: Neuron distributions over the layers of Mistral-Nemo-12B for all 12 languages. Asterisks indicate languages with non-Latin scripts.
Lang. # Neurons # Trainable %
@@ B.1 Language Subnetworks
Figure 9: Neuron distributions over the layers of Aya-Expanse-8B for all 12 languages. Asterisks indicate languages with non-Latin scripts.
Lang. # Neurons # Trainable %
Afrikaans 2091 25.7M 0.32
Welsh 5820 71.5M 0.88
Estonian 3628 44.6M 0.55
Icelandic 4303 52.9M 0.65
Georgian 2432 29.9M 0.37
Lithuanian 3443 42.3M 0.52
Latvian 3536 43.5M 0.54
Macedonian 2439 30.0M 0.37
Maltese 3201 39.3M 0.49
Nepali 3870 47.6M 0.59
Slovak 1924 23.6M 0.29
Slovenian 2538 31.2M 0.39
Figure 10: Neuron and parameter counts for different language subnetworks, and percentages of the model being trained within Aya-Expanse-8B.
B.2 Language Neuron Overlap
Figure 11 shows pairwise overlap between language-specific subnetworks, grouped by language family. Across models, related languages consistently exhibit higher overlap than unrelated ones, indicating partial sharing of language-relevant components rather than strictly disjoint representations.
(a) Llama-3.1-8B
(b) Mistral-Nemo-12B
(c) Aya-Expanse-8B
Figure 11: Language neuron overlap for all 12 languages across models.
B.3 Language Neuron Distributions
Below, we provide per-language distributions of identified language-specific neurons across layers for each model.
(a) Llama-3.1-8B
(b) Mistral-Nemo-12B
(c) Aya-Expanse-8B
Figure 12: Language neuron distributions for all 12 languages across models.
Appendix C Training Dynamics
This section presents validation loss curves for all 12 target languages, illustrating training dynamics across fine-tuning strategies.
(a) Afrikaans
(b) Welsh
(c) Estonian
(d) Icelandic
(e) Georgian
(f) Lithuanian
(g) Latvian
(h) Macedonian
(i) Maltese
(j) Nepali
(k) Slovak
(l) Slovenian
Figure 13: Validation loss curves for Llama-3.1-8B across languages.
Appendix D Downstream Results
In this appendix, we provide detailed per-language and per-task performance results for all models, complementing the averaged results reported in the main text.
D.1 Llama-3.1-8B
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@ Appendix D Downstream Results
arc challenge 70.90 25.75 35.45 64.55 73.91 62.21 71.24
arc easy 85.79 26.14 53.33 79.12 85.96 81.75 85.26
hellaswag 40.38 14.68 26.13 35.22 38.26 36.04 38.72
piqa 71.87 39.17 50.00 73.07 74.05 72.03 71.71
winogrande 51.38 49.17 50.43 50.91 52.49 47.04 51.46
mmlu 58.07 27.63 37.30 51.60 57.61 54.74 56.43
Avg. 63.07 30.42 42.11 59.08 63.71 58.97 62.47
Table 3: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 16.05 22.74 56.52 69.57 54.52 68.23
arc easy 85.79 16.84 30.70 74.56 85.61 76.14 84.04
hellaswag 40.38 12.31 25.27 29.93 38.66 30.88 40.79
piqa 71.87 3.97 49.46 58.98 73.72 66.97 68.39
winogrande 51.62 20.28 49.57 49.17 50.20 51.46 50.51
mmlu 58.07 22.27 25.67 45.79 58.26 48.14 56.69
Avg. 63.10 15.29 33.90 52.49 62.67 54.68 61.44
Table 4: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 22.07 36.45 58.86 71.57 63.55 72.58
arc easy 85.79 21.40 55.09 77.89 84.91 81.58 84.74
hellaswag 40.38 13.92 22.14 36.11 39.84 37.06 41.82
piqa 71.87 40.10 54.35 59.63 75.14 72.96 75.24
winogrande 51.38 41.83 50.20 51.14 52.80 52.01 53.83
mmlu 58.07 20.64 33.90 48.66 58.59 52.84 57.35
Avg. 63.07 26.66 42.02 55.38 63.81 60.00 64.26
Table 5: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 21.74 37.79 56.19 68.23 59.20 74.25
arc easy 85.79 23.33 55.96 74.39 83.68 76.84 85.61
hellaswag 40.38 8.29 25.24 32.65 40.81 27.75 44.67
piqa 71.87 26.55 54.79 69.48 76.01 61.48 77.31
winogrande 51.38 40.02 50.28 52.01 51.93 52.17 54.54
mmlu 58.07 17.18 35.86 46.77 57.09 45.40 58.92
Avg. 63.07 22.85 43.32 55.25 62.96 53.81 65.88
Table 6: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 9.03 10.70 71.57 70.90 63.55 69.90
arc easy 85.79 10.18 15.44 85.26 85.96 82.98 84.56
hellaswag 40.38 1.24 6.33 42.39 42.30 33.13 41.56
piqa 71.87 12.13 8.81 73.67 72.69 64.53 71.16
winogrande 51.38 3.95 37.57 53.12 52.64 50.91 51.38
mmlu 58.07 4.51 18.42 58.00 58.00 55.26 56.69
Avg. 63.07 6.84 16.21 64.00 63.75 58.39 62.54
Table 7: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 20.74 37.79 54.52 69.90 67.89 70.23
arc easy 85.79 27.89 59.12 74.04 84.56 80.53 85.26
hellaswag 40.38 20.49 29.02 31.45 41.50 38.14 41.37
piqa 71.87 39.28 51.63 57.56 74.37 69.15 74.21
winogrande 51.38 51.07 49.96 50.43 51.62 50.36 53.35
mmlu 58.07 27.24 36.64 45.85 58.20 52.19 57.02
Avg. 63.07 31.12 44.03 52.31 63.36 59.71 63.58
Table 8: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 20.07 25.42 53.51 71.24 68.23 69.23
arc easy 85.79 27.54 36.67 70.70 84.21 81.58 84.56
hellaswag 40.38 14.48 17.34 31.23 40.71 35.03 41.73
piqa 71.87 39.50 43.69 55.93 75.46 71.82 74.59
winogrande 51.62 47.43 46.80 50.04 52.57 50.43 55.17
mmlu 58.07 24.89 30.70 45.59 58.07 53.43 59.63
Avg. 63.10 28.98 33.44 51.17 63.71 60.09 64.15
Table 9: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 31.10 43.48 69.90 72.24 66.22 69.23
arc easy 85.79 39.82 64.39 84.39 85.61 82.11 85.61
hellaswag 40.38 26.41 25.06 37.91 41.80 32.57 36.43
piqa 71.87 49.67 53.26 73.72 75.73 68.44 67.36
winogrande 51.62 49.41 49.25 52.96 52.25 51.78 52.09
mmlu 58.07 30.96 39.65 56.69 58.39 54.67 57.48
Avg. 63.10 37.90 45.85 62.60 64.34 59.30 61.37
Table 10: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 22.07 38.13 49.16 69.57 62.21 71.57
arc easy 85.79 21.23 56.84 72.46 83.51 82.11 83.86
hellaswag 40.38 9.70 26.72 31.45 36.30 35.07 39.00
piqa 71.87 42.11 50.87 61.97 70.73 72.85 73.34
winogrande 51.54 51.22 49.17 49.96 50.75 53.67 50.91
mmlu 58.07 26.26 32.66 45.00 55.91 52.58 57.87
Avg. 63.09 28.77 42.40 51.67 61.13 59.75 62.76
Table 11: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 12.71 35.45 65.89 69.90 67.56 71.57
arc easy 85.79 19.47 49.12 81.75 82.46 82.98 84.39
hellaswag 40.38 16.38 14.32 37.46 36.73 36.68 36.31
piqa 71.87 36.83 51.25 62.35 74.59 73.83 71.11
winogrande 51.38 13.58 47.04 50.12 52.41 54.06 50.99
mmlu 58.07 17.50 33.18 54.08 57.67 54.08 57.74
Avg. 63.07 19.41 38.39 58.61 62.29 61.53 62.02
Table 12: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 21.74 36.45 61.20 71.91 62.21 71.57
arc easy 85.79 23.86 51.05 78.25 86.49 78.60 85.44
hellaswag 40.38 15.04 27.87 40.04 40.76 36.74 43.68
piqa 71.87 43.25 54.30 59.25 72.91 70.29 71.76
winogrande 51.54 38.91 50.43 51.07 53.04 50.91 51.70
mmlu 58.07 21.36 35.99 52.25 58.20 51.53 57.67
Avg. 63.09 27.36 42.68 57.01 63.88 58.38 63.64
Table 13: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
@@
arc challenge 70.90 21.74 34.78 53.51 67.89 67.56 67.89
arc easy 85.79 28.60 36.49 71.23 84.39 82.46 83.51
hellaswag 40.38 8.79 24.92 34.08 39.98 39.72 36.48
piqa 71.87 45.65 56.15 70.13 74.43 72.03 75.08
winogrande 51.46 49.88 50.67 49.64 52.17 50.43 52.25
mmlu 58.07 27.43 29.13 43.31 57.94 55.06 56.69
Avg. 63.08 30.35 38.69 53.65 62.80 61.21 61.98
Table 14: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian for Llama-3.1-8B.
D.2 Mistral-Nemo-12B
@@
Table 15: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans for Mistral-Nemo-12B.
@@
Table 16: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh for Mistral-Nemo-12B.
@@
Table 17: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian for Mistral-Nemo-12B.
@@
Table 18: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic for Mistral-Nemo-12B.
@@
Table 19: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian for Mistral-Nemo-12B.
@@
Table 20: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian for Mistral-Nemo-12B.
@@
Table 21: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian for Mistral-Nemo-12B.
@@
Table 22: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian for Mistral-Nemo-12B.
@@
Table 23: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese for Mistral-Nemo-12B.
@@
Table 24: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali for Mistral-Nemo-12B.
@@
Table 25: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak for Mistral-Nemo-12B.
@@
Table 26: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian for Mistral-Nemo-12B.
D.3 Aya-Expanse-8B
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE af 76.89 81.00 81.56
SIB200 af 42.16 49.51 64.22
[…]
Avg. 36.84 42.70 44.25
General Performance
arc challenge 77.26 78.26 75.59
arc easy 90.00 88.95 90.53
hellaswag 56.15 43.21 40.11
piqa 77.48 78.35 79.98
winogrande 55.72 54.22 53.43
mmlu 88.11 83.08 85.76
Avg. 74.12 71.01 70.90
Table 27: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
SIB200 cy 32.35 67.65 58.33
[…]
Avg. 17.13 29.64 28.63
General Performance
arc challenge 77.26 72.24 76.25
arc easy 90.00 87.72 88.25
hellaswag 56.15 38.89 55.82
piqa 77.48 77.64 81.23
winogrande 55.72 56.12 56.04
mmlu 88.11 77.14 83.80
Avg. 74.12 68.29 73.56
Table 28: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE et 54.00 71.78 68.78
SIB200 et 39.71 50.49 66.67
[…]
Avg. 20.93 33.36 33.61
General Performance
arc challenge 77.26 73.91 73.24
arc easy 90.00 89.12 87.54
hellaswag 56.15 49.79 44.84
piqa 77.48 76.17 80.47
winogrande 55.72 55.09 57.30
mmlu 88.11 82.04 84.78
Avg. 74.12 71.02 71.36
Table 29: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE is 53.22 70.00 72.78
SIB200 is 33.33 59.80 62.75
[…]
Avg. 21.07 33.50 34.06
General Performance
arc challenge 77.26 75.59 77.59
arc easy 90.00 88.77 89.82
hellaswag 56.15 38.54 45.24
piqa 77.48 76.55 79.60
winogrande 55.72 54.22 53.75
mmlu 88.11 85.56 87.79
Avg. 74.12 69.87 72.30
Table 30: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE ka 53.56 58.44 60.44
SIB200 ka 48.04 68.63 64.71
[…]
Avg. 21.74 27.08 26.90
General Performance
arc challenge 77.26 73.58 77.93
arc easy 90.00 88.07 89.82
hellaswag 56.15 44.69 52.88
piqa 77.48 81.56 81.56
winogrande 55.72 52.96 54.85
mmlu 88.11 84.00 87.07
Avg. 74.12 70.81 74.02
Table 31: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE lt 70.22 74.67 76.22
SIB200 lt 45.59 43.63 46.08
[…]
Avg. 28.56 30.45 33.32
General Performance
arc challenge 77.26 73.91 74.92
arc easy 90.00 86.84 87.54
hellaswag 56.15 46.50 48.36
piqa 77.48 78.78 80.69
winogrande 55.72 53.35 55.88
mmlu 88.11 82.95 86.28
Avg. 74.12 70.39 72.28
Table 32: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE lv 56.56 77.33 75.78
SIB200 lv 42.16 62.75 67.16
[…]
Avg. 23.21 38.17 38.03
General Performance
arc challenge 77.26 69.23 77.93
arc easy 90.00 86.14 89.47
hellaswag 56.15 42.04 52.23
piqa 77.48 77.26 81.88
winogrande 55.72 55.09 54.85
mmlu 88.11 81.71 87.20
Avg. 74.12 68.58 73.93
Table 33: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE mk 67.33 77.22 78.67
SIB200 mk 43.14 66.67 70.10
[…]
Avg. 28.24 41.03 42.24
General Performance
arc challenge 77.26 75.59 78.60
arc easy 90.00 87.89 90.00
hellaswag 56.15 48.53 43.78
piqa 77.48 78.18 82.05
winogrande 55.72 55.56 54.30
mmlu 88.11 84.39 88.18
Avg. 74.12 71.69 72.82
Table 34: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE mt 43.56 74.78 74.56
SIB200 mt 45.59 59.80 68.63
[…]
Avg. 22.97 43.76 44.47
General Performance
arc challenge 77.26 76.92 74.58
arc easy 90.00 87.89 87.89
hellaswag 56.15 39.55 42.98
piqa 77.48 79.27 79.76
winogrande 55.72 53.28 55.09
mmlu 88.11 84.32 85.50
Avg. 74.12 70.21 70.97
Table 35: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE ne 45.89 63.11 66.00
SIB200 ne 47.55 65.20 69.12
[…]
Avg. 23.96 34.03 34.87
General Performance
arc challenge 77.26 73.24 79.26
arc easy 90.00 87.72 90.88
hellaswag 56.15 41.11 49.52
piqa 77.48 81.28 82.64
winogrande 55.72 53.99 54.85
mmlu 88.11 85.50 88.96
Avg. 74.12 70.47 74.35
Table 36: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE sk 80.56 79.33 81.00
SIB200 sk 52.94 42.16 54.90
[…]
Avg. 36.92 36.28 38.99
General Performance
arc challenge 77.26 77.59 74.92
arc easy 90.00 87.72 88.77
hellaswag 56.15 52.13 48.54
piqa 77.48 79.38 79.16
winogrande 55.72 53.99 54.70
mmlu 88.11 83.87 86.02
Avg. 74.12 72.45 72.02
Table 37: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE sl 67.22 78.56 80.22
SIB200 sl 40.69 49.51 54.90
[…]
Avg. 27.05 34.50 36.00
General Performance
arc challenge 77.26 69.23 78.60
arc easy 90.00 85.79 89.82
hellaswag 56.15 47.88 51.96
piqa 77.48 76.61 79.87
winogrande 55.72 54.93 57.46
mmlu 88.11 80.54 86.81
Avg. 74.12 69.16 74.09
Table 38: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian for Aya-Expanse-8B.
Appendix E Generation Prompts
E.1 Task Prompts